OpenStudy (anonymous):

Help with integration problem. Need to integrate 1 / ( 1 + x^(3/2)) dx Is there a way to do this without using partial fractions?

OpenStudy (anonymous):

\[\int\limits_{}^{} 1/(1 + \sqrt[3]{x}) dx\]

OpenStudy (amistre64):

maybe z^2 = x^(1/3) ?? and see if that gets you a good set up, might not tho

OpenStudy (anonymous):

hmm, let me try...

OpenStudy (amistre64):

otherwise you might try to see if wolfram has a good step down of it

OpenStudy (anonymous):

tried a bit but am getting lost in circles. Wolfram uses partial fractions...let me look at your link...

OpenStudy (amistre64):

if we let: z^6 = x 6z^5 dz = dx \[\int\frac{6z^4}{1+z^2}dz\] and integrate by parts?

OpenStudy (anonymous):

aye, that might work.

OpenStudy (amistre64):

it might, but then it might not :)

OpenStudy (amistre64):

you could then take out the constant 6; divide out the rest and do a sum of the dividends

OpenStudy (amistre64):

and that should be a 6z^5 .... typoed it

OpenStudy (anonymous):

How did you get z^6? Doesnt z = x^3/2 turn into z^2/3 = x ?

OpenStudy (amistre64):

\[z^6=x;\ z^{6/3}=z^2\] \[6\left(\frac{z^4}{4}\frac{z^2}{2}+\frac{1}{2}\int\frac{2z}{z^2+1}\right)\] \[6\left(\frac{z^4}{4}\frac{z^2}{2}+\frac{tan^{-1}(z)}{2}\right)\]

OpenStudy (amistre64):

\[6\left(\frac{z^4}{4}-\frac{z^2}{2}+\frac{tan^{-1}(z)}{2}\right)\]

OpenStudy (amistre64):

you simply change the variable to z^6 = x to accomodate for the rest

OpenStudy (amistre64):

ive seen it in the books before, but i havent had much practice at it

OpenStudy (amistre64):

and it looks prime for this set up tho

OpenStudy (amistre64):

z^3 - z + z/(z^2+1) ------------------ z^2 +1 | z^5 -z^3 +z

OpenStudy (amistre64):

i see a fauxpaux in the integral i did ..... urgh

OpenStudy (anonymous):

this one is a tough one...making my head hurt! :)

OpenStudy (amistre64):

\[6\left(\frac{z^4}{4}-\frac{z^2}{2}+\frac{1}{2}\int\frac{2z}{z^2+1}\right)\] \[6\left(\frac{z^4}{4}-\frac{z^2}{2}+\frac{ln|2x+1|}{2}\right)\]

OpenStudy (amistre64):

and of course 2x menas x^2; typing and mathing just dont mix

OpenStudy (amistre64):

im sure it doesnt work in the end, but its a good line to try to figure out :)

OpenStudy (anonymous):

I'm a little lost a step here, let me study this for a bit. Mind watching this thread when its updated by me?

OpenStudy (amistre64):

if im around or maybe tomorrow